OpenStudy (anonymous):

write the standard form of the equation line through the given point. through:(-3,2) and(0,-1)

OpenStudy (anonymous):

help me formular

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

m=-3/3=-1 choose any x1 and y1

OpenStudy (anonymous):

ok thank you